OpenStudy (anonymous):

Evaluate the determint of the matrix |1 2| |-2 -1| this is one matrix

OpenStudy (anonymous):

|A|=3

OpenStudy (anonymous):

may you explain ?

OpenStudy (anonymous):

for 2 by 2 matrix

OpenStudy (anonymous):

\[\left[\begin{matrix}1 & 2 \\ -2 & -1\end{matrix}\right]=1*(-1)-(-2)*2=-1+4 =3\]

OpenStudy (anonymous):

thanks can we do 1 more?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

but after the medal :p

OpenStudy (anonymous):

know what post new question this way i get one more hehe

OpenStudy (anonymous):

oh sorry ...and i will